OpenStudy (greencat):

@sammixboo and your pie love: I have a new favorite number for you... and it's kinda lame :/

sammixboo (sammixboo):

:o?

sammixboo (sammixboo):

Tell me this number

OpenStudy (greencat):

\[\sqrt{-64}\times \pi\]

OpenStudy (greencat):

Do the math :D

OpenStudy (greencat):

@sammixboo It would be \(i8\pi\).

OpenStudy (greencat):

;-; #ignored

OpenStudy (greencat):

Now I know what Yana and amq were talking about.

Nnesha (nnesha):

\[\huge\rm\sqrt{-64} \times \pi\] there is negative sign under the square root so you have to separate that square root of negative one and 64

Nnesha (nnesha):

\[\large\rm \sqrt{-1} = i ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ i^2 = -1\] so